PROBLEM:
DAVE IS FLYING FROM PHILADELPHIA TO CHICAGO. ON THE FLIGHT, FLYING AGAINST THE WIND, THE 780 MILE TRIP TAKES 2 HOURS. THE RETURN TRIP, WITH THE WIND TAKES 1.5 HOURS. FIND THE SPEED OF THE WIND AND FIND THE SPEED OF THE PLANE IN STILL AIR.
PLEASE HELP I DONT EVEN KNOW WHERE TO BEGIN THANKS!!

Answer by checkley77(12844)

780/2=390 MPH AGAINST THE WIND
780/1.5=520 MPH WITH THE WIND
(520-390)/2=130/2=65 MPH FOR THE WIND SPED.
PROOF:
520-65=390+65
455=455 THIS IS THE SPEED OF THE PLANE IN STILL AIR.